THE GROUND STATE ROTATIONAL SPECTRUM OF SO_{2}F_{2}.

\underline{M. ROTGER}, V. BOUDON and M. LOËTE, Laboratoire de Physique de l’Université de Bourgogne, CNRS UMR 5027, 9, Avenue Alain Savary, B.P. 47 870, F-21078 DIJON Cedex, France; L. MARGULÈS and J. DEMAISON, Laboratoire de Physique, Atomes et Molécules, CNRS UMR 8523, Université de Lille I, Bât. P5, 59655 Villeneuve d’Ascq Cedex, France; H. MÄDER, Institut für Physikalische Chemie, Universität Kiel, Olshausenstr. 40, D-24098 KIEL Germany; G. WIN-NEWISSER and H. S. P. MÜLLER, I. Physikalisches Institut, Universität zu Köln, D-50937 KÖLN Germany.

The analysis of the ground state rotational spectrum of SO_{2}F_{2}^{a} has been performed with the Watson’s Hamiltonian up to sextic terms but shows some limits due to the A and S reductions. Since SO_{2}F_{2} is a quasi-spherical top, it can also be regarded as derived from an hypothetical XY_{4} molecule. Thus we have developped a new tensorial formalism in the O(3) ⊃ T_{d} ⊃ C_{2v} group chain^{b}. We test it on the ground state of this molecule using the same experimental data^{c} (0–1 THz region, J up to 99). Both fits are comparable even if the formalisms are slightly different. This talk intends to establish a link between the classical approach and the tensorial formalism. In particular, our tensorial parameters at a given order of the development are related to the usual ones. Programs for spectrum simulation and fit using these methods are named C_{2v}TDS. They are freely available at the URL:

            http://www.u-bourgogne.fr/LPUB/c2vTDS.html

aK. Sarka, J. Demaison, L. Margulès, I. Merke, N. Heineking, H. Bürger and H. Ruland, J. Mol. Spectrosc., 200, 55-64, (2000).

bM. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc., 216, 297-307, (2002).

cM. Rotger, V. Boudon, M. Loëte, L. Margulès, J. Demaison, H. Mäder, G. Winnewisser and H.S.P. Müller, J. Mol. Spectrosc., in preparation, (2003).